OpenStudy (anonymous):

How would i sketch y = Ix + 1I

OpenStudy (anonymous):

Start by drawing a graph with x values going from -5 to 5 and y values going from 0 upwards. Then put values in for x between -5 and 5. so lets start with x=0. y(0)=abs(0+1)=1 where abs means absolute. Absolute value means y can not be negative for this function(which is why we plot the graph going from 0 up for 'y' values). eg. x=-5 y(-5)=abs(-5+1)=abs(-4)=4.

OpenStudy (anonymous):

do u knw the sketching of graph y =modx

OpenStudy (anonymous):

mod?

OpenStudy (anonymous):

draw the graph of y=|x| first and then shift the vertex to -1,0 graph of y=|x| is V sahped symmetrical about y-axis and inclined at 45 wrt x-axis

OpenStudy (anonymous):

if u know the graph of modulus x the apply curve transformation by shift the vertex to -1,0

OpenStudy (zzr0ck3r):

lol omg mod to graph abs(x)???????? ....graph abs(x) then move it to the left 1

OpenStudy (zzr0ck3r):



OpenStudy (anonymous):